LUCI2/AO Commissioning 20150510 UT

Observers: DMiller, GBrusa, JHill, SW Group (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support: DMiller (LBTO)
Software Support: (Tucson)

Summary

Details

Afternoon test with Doug and Guido

testing SX AdSec after instalation at 3:45pm on May 21

we found that right after turning on several actuators were giving 0.2m\x85

41, 3,2,368,27,58,60 (MIR)

we stopped the secondary and restarted

removed actuator 2 (MIR but also DSP)

used fsm_set_flat with default and found there was a problem with HO offloading

set the astigmatism LUT to all zeros and try again

we finally got to acquire a new flat and we have actuator 58 with a bit higher force (perhaps it should go to icc list?)

we tested the mirror down to 32 with shell set and it was ok

after resting the shell we found 541 DSP went to 0.25m

The following links have been created:

ADOPT_ROOT: /home/aoacct/AO/current --> /home/aoacct/AO/FLAO2-20120129 ADOPT_LOG: /local/aolog/current --> /local/aolog/FLAO2-20120129

[AOeng@adsecsx src4]$ adsc_start Configuration directory: /home/aoacct/AO/current/conf/adsec/current/processConf Side: L Subsystem: ADSEC MsgD-RTDB is active. Please execute a stop command before.


%USERSIG{DougMiller - 2015-05-22}%
Topic revision: r2 - 22 May 2015, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback